what is the additive inverse (opposite) of -5
Found 2 solutions by Boreal, Shin123:
Answer by Boreal(15235)

+5. The additive inverse is the number which when added to the given number, equals 0.

Answer by Shin123(626)

-5+__=0 The additive inverse is that if you add the additive inverse to the number, you will get 0.
0-(-5)=__ 0+5=__ 0+5=5
The correct answer is 5.